Please help me solve this system of equations:
7x+5y+z=0
-x+3y+2z=16
x-6y-z=-18
Thank you!

Found 2 solutions by palanisamy, vksarvepalli:
Answer by palanisamy(496)

The given equations are,
7x+5y+z=0 ...(1)
-x+3y+2z=16 ...(2)
x-6y-z=-18 ...(3)
(1)*2=> 14x+10y+2z = 0 ...(4)
(2)-(4)=> -15x-7y = 16 ...(5)
(1)+(3)=> 8x-y = -18 ...(6)
(6)*7=> 56x-7y = -126 ...(7)
(5)-(7)=> -71x = 142
x = 142/(-71)
x = -2
(6)=> 8*(-2)-y = -18
-16-y = -18
-y = -18+16
-y = -2
y = 2
(1)=> 7*(-2)+5*(2)+z = 0
-14+10+z= 0
-4+z = 0
z = 4
Therefore the solution is, x= -2, y=2 and z= 4

Answer by vksarvepalli(154)

7x+5y+z=0 ------------ 1
-x+3y+2z=16 ------------ 2
x-6y-z=-18 ------------ 3
adding Eq.s 1 & 3
we get 8x-y=-18 --- 4
multiply eq.3 by 2 and add to eq.2
we get x-9y=-20 --- 5
multiply eq.5 by 8 and subtract it from eq.4
we get 71y=142
so y=2
from eq. 5 x-18=-20
so x=-2
from eq.1 -14+10+z=0
so z=4
x=-2, y=2, z=4